An artificial satellite circling the Earth completes each orbit in 110 minutes. a.) Find the altitude of the satellite. b.) What is the value of g at the locatioin ofthis satellite?
Answer #1
Period of satellite T = 110 min= 110 x 60 s=6600 sWe know T = 2π √[ r 3 / GM ]r 3 / GM = (T/2π) 2r 3 = GM (T/2π) 2Where G = Gravitational constant= 6.67 x 10 -11 N m 2 / kg2M = Mass of earth= 5.98x 10 24 kgSubstitute the values we get r 3 = 4.401 x 10 20 m 3r = 7.606 x 10 6 m∴Altitide of the satellite h = r -RWhere R = radius of earth= 6.38 x 10 6 mSo, h = 1.226 x 10 6 m(b). The value of g at the locatioin of this satellite g ' = g [ R /r ] 2= g [0.7036] where g = 9.8 m /s 2= 6.89 m / s 2
answered by: unknown!